OpenStudy (anonymous):

find the real and imaginary solutions to the following equation: 7(l4x-1l)-10=(l4x-1l)^2... the lines in the parenthesis are absolute value signs

OpenStudy (anonymous):

Let W=4x-11 \(\large \rightarrow 7|W|-10=|W|^2\). It doesn't matter for the W^2 if W is positive or negative, so there are only two possibilities: \(\large 7W-10=W^2\), or \(\large -7W-10=W^2\). Solve each of those for W, then substitute 4x-11 back in to find x.

OpenStudy (anonymous):

i am confused on how to solve for w here

OpenStudy (anonymous):

They are quadratic equations. You can put them in the form aW^2+bW+c=0, then factor or use quadratic formula.

OpenStudy (anonymous):

can you please start me off im am new with this and still confused. thanks

OpenStudy (anonymous):

Starting with \(7W−10=W^2\), rearrange it to be \(W^2-7W+10=0\). This can be factored to \((W-2)(W-5)=0\). You can the sub back in the W=4x-11 to get 4x-11=2 and 4x-11=5.

OpenStudy (anonymous):

but when you bring w^2 to the other side wouldnt it be negative

OpenStudy (anonymous):

I didn't bring W^2 to the other side, I brought everything else over to be with W^2.

OpenStudy (anonymous):

nevermind i understood that part

OpenStudy (anonymous):

but it is not 4x-11... it is 4x-1

OpenStudy (anonymous):

LOL, sorry, the || were making my eyes cross.

OpenStudy (anonymous):

yea sorry about that

OpenStudy (anonymous):

It still works the same way. I'm redoing it to get the correct answers now.

OpenStudy (anonymous):

I found four real solutions, but no imaginary ones.

OpenStudy (anonymous):

i only found 2 real solutions... x=3/4 and x=1.5

OpenStudy (anonymous):

Ok, those are the correct ones from the first equation, now you need to do the same procedure on \(W^2+7W+10=0\).

OpenStudy (anonymous):

what were your final solutions?

OpenStudy (anonymous):

x \(\epsilon\) {3/4, 3/2, -1/4, and one other one that I'll let you get on your own}

OpenStudy (anonymous):

i do not understand how to get -1/4

OpenStudy (anonymous):

\[ x=a+ib \] \[ 7(|4x-1|)-10=|4x-1|^2 \] \[ 7\sqrt{(4a-1)^2+(4b)^2}-10=\sqrt{(4a-1)^2+(4b)^2}^2 \] \[ 7\sqrt{(4a-1)^2+(4b)^2}=16a^2-8a+1+16b^2+10 \]

OpenStudy (anonymous):

i need to use u-substitution

OpenStudy (anonymous):

CliffSedge may you please help me find the last two real solutions?

OpenStudy (anonymous):

Same as before. \(W^2+7W+10=0=(W+2)(W+5)\) \(\rightarrow 4x-1=-2\), and \(4x-1=-5\) I'm going through henpen's equation to see if I can find any imaginary solutions.

OpenStudy (anonymous):

my final answers came out to be x=3/4, 1.5, -1, -1/4

OpenStudy (anonymous):

is this right?

OpenStudy (anonymous):

That's what I got too. Well done.

OpenStudy (anonymous):

Were you able to make any sense of what henpen posted above?